OpenStudy (anonymous):

How would you factor x^2-x-12

OpenStudy (anonymous):

I got something like (x+1)(x-12) Something in that range. Not too sure though.

OpenStudy (anonymous):

what two numbers when added = -1 and when multiplied = -12?

OpenStudy (anonymous):

(x+1)(x-12) = x^2 + x - 12x - 12 So you had the x^2 and -12 parts ok, but your factors don't produce the "-x" in the middle. If you (x + a) and (x - b) where a * b = -12 but a - b = -1, then you could factor like that...

OpenStudy (anonymous):

think of other pairs of numbers besides 12 and 1 that multiply to make 12... try them as factors...

OpenStudy (anonymous):

Hmm.. I see where you're coming at. Give me a bit.

OpenStudy (anonymous):

So that would be (x+3)(x-4)

OpenStudy (anonymous):

:-)

OpenStudy (anonymous):

._.?

OpenStudy (anonymous):

your answer is correct.

OpenStudy (anonymous):

good work @Jpigott :)

OpenStudy (anonymous):

Great! The answer to my question is: \[(-\infty,-4)\upsilon(-4,3)\upsilon(3,\infty)\]

OpenStudy (anonymous):

I had to find the domain for it. Thanks :)